Can you please help me solve the following problem.
Tomorrow,the number of candles on Claire's birthday cake will be one less than three times two-fifths of the number five years ago.How old is Claire today?

Answer by stanbon(75887)

Tomorrow,the number of candles on Claire's birthday cake will be one less than three times two-fifths of the number five years ago.
How old is Claire today?
------
Equation:
x = 3(2/5)(x-5)-1
---
x+1 = (6/5)(x-5)
----
5x+5 = 6x-30
x = 35 (Claire's age today)
================================
Cheers,
Stan H.
===============